Diffusion Morning Report 09/13/07

Ted Berg tberg at stanford.edu
Thu Sep 13 07:11:45 PDT 2007


Good Morning All,
    Looking Ok this Morning..Many tools need testing. Pocl3 and BBr3 
still under observation. Ebeam no beam current.

  	  	  	  	  	 
  	Date: 	9/13/07 	THICKNESS-Change at 	  	 
Tube Name/Number 	Status 	  	  	  	 
Tylan 1 	down  	door close issue 	  	  	 
Tylan 2 	up 	  	  	  	 
Tylan 3 	up 	  	  	  	 
Tylan 4 	up 	  	  	  	 
Tylan 5-BBr3 	yellow 	observation 	  	  	 
Tylan 6-POCl3 	yellow 	smoking 	  	  	 
Tylan 7-TEOS2 	yellow 	needs testing 	2.86/30 	  	 
Tylan 8-SiGe 	yellow 	needs testing 	21.64/30 	  	 
Tylan 9-Poly 	up 	  	24.58/30 	  	 
Tylan 10-Nitride 	up 	  	.3/30 	  	 
Tylan 11-BPSG-LTO 	up 	  	8.36/30 	  	 
Tylan 12-FGA 	up 	  	  	  	 
  	  	  	  	  	 
Thermco 1 	up 	  	  	  	 
Thermco 2 	up 	  	  	  	 
Thermco 3 	future 	  	  	  	 
Thermco4 	future 	  	  	  	 
Thermco Poly1 	yellow 	testing 	  	  	 
Thermco Poly2 	future 	needs testing 	
	
	
Thermco Nitride1 	up 	  	  	  	 
Thermco Nitride2 	future 	
	  	  	 
Burn Box 	Up 	  	   	  	 
Ebeam 	down  	no beam current 	  	  	 
ZYGO 	up 	  	  	  	 
FGA2 	up 	  	  	  	 
SiH4 for furnaces 	up 	A on line 510 psi 	 changeover working well  	  	 


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20070913/ef288143/attachment.html>


More information about the diffusion mailing list